Calculate the pH of each of the following strong acid solutions.



(a) 0.00890 M HI

pH =



(b) 0.901 g of HClO4 in 28.0 L of solution

pH =



(c) 17.0 mL of 3.70 M HI diluted to 4.70 L

pH =



(d) a mixture formed by adding 12.0 mL of 0.00655 M HI to 72.0 mL of 0.000920 M HClO4

pH =
Answer #1

H+ + I) 0.00890 MHI CHI CH+] = 0.00890M pH = a log CHT] - log (0.00890) 2.05 M= 0.901 - 28.04 = 3.9x10 100.46. pH = -log (3.2